OpenStudy (anonymous):

Analyze each one in the following way: find the domain and range, x and y intercept(s), horizontal asymptote(s), and vertical asymptote(s): f(x) = 2 / x^2 - 2x - 3

OpenStudy (anonymous):

@timo86m

OpenStudy (anonymous):

you mean 2 / (x^2 - 2x - 3) Range is all y values that work. domain is all x values that work. Here is a graph of it https://www.google.com/search?q=2+%2F+(x%5E2+-+2x+-+3)&oq=2+%2F+(x%5E2+-+2x+-+3)&aqs=chrome..69i57j0l3.658j0&sourceid=chrome&ie=UTF-8 we can see all y values are in the range except for 0. Which is also the horizontal asymptote. ---- As for domain those are all the x values that are defined and work for the function You can interpret this as no division by 0. SO take the denominator portion of the function and set it to 0 (x^2 - 2x - 3)=0 -1, 3 are you answers when you solve for it. So DOmain is all real numbers excep -1 and 3. ALso it is your V asimptotes Range is all rea numbers except 0. Also it is H asympote.

OpenStudy (anonymous):

That takes care of domain and range and your asymptotes :D

OpenStudy (anonymous):

As for x intercepts you set f(x)=0 and solve. In other words 2 / (x^2 - 2x - 3)=0 THis is a little harder try it first then let me know if you need help as for y intercept you go f(0) in other words plug in 0 to every x in f(x) as so 2 / (0^2 - 2*0 - 3)=y -.66666=y